Vatic Capital.  Cachita Haynes works as a currency speculator for Vatic Capital of Los Angeles. Her latest speculative position is to profit from her expectation that the U.S. dollar will rise significantly against the Japanese yen. The current spot rate is ​¥121.00/$. She must choose between the following 90​-dayoptions on the Japanese​ yen:

Option

Strike Price

Premium

Put on yen

yen¥125/$

​$0.00003​/yen¥

Call on yen

yen¥125​/$

​$0.00046​/yen

a. Should Cachita buy a put on yen or a call on​ yen?

b. What is​ Cachita's breakeven price on the option purchased in part a​?

c. Using your answer from part a​, what is​ Cachita's gross profit and net profit​ (including premium) if the spot rate at the end of 90 days is ​¥140.00​/$?
Answer #1

a. Since her view is that the Yen would depreciate, she should buy a call on USD/JPY i.e. a Put on Yen

b. The current spot exchange rate is 1/121 USD per yen = 0.008264

Her strike price is 1/125 = 0.008

If she buys a put on the yen at a cost of USD 0.00003 per yen, her break even would be 0.008 - 0.00003 = 0.00797 i.e. 125.4705 in terms of Yen per USD

c. If after 90 days, USD/JPY is at 140 i.e. JPY/USD is at 1/140 = 0.007143, then her gross profit would be 0.008 - 0.007143 = 0.000857 USD per yen

Net profit would be 0.000857 - 0.00003 = 0.000827 USD per yen